uppose the amounts presented here are basic financial information (in millions) from the 2020 annual reports of Nike and adidas.

Nike

adidas

Sales revenue $18,954.3 $10,251
Allowance for doubtful accounts, beginning 76.8 116
Allowance for doubtful accounts, ending 107.1 131
Accounts receivable balance (gross), beginning 2,886.9 1,754
Accounts receivable balance (gross), ending 2,955.0 1,553

Collapse question part

(a1)

Calculate the accounts receivable turnover for both companies. (Round answers to 1 decimal place, e.g. 12.5.)

Nike

adidas

Accounts receivable turnover

times

times
Answer #1

SOLUTION

Nike Adidas
Accounts receivable balance (gross), ending 2,955.0 1,553
Less: Allowance for doubtful accounts, ending 107.1 131
Accounts receivable ending balance (Net) (A) 2,847.9 1,422
Accounts receivable balance (gross), beginning 2,886.9 1,754
Less: Allowance for doubtful accounts, beginning 76.8 116
Accounts receivable beginning balance (Net) (B) 2,810.1 1,638
Average accounts receivable [(A+B)/2] 2,829 1,530

Accounts receivable turnover = Sales / Average accounts receivable

Nike-

Accounts receivable turnover = 18,954.3 / 2,829 = 6.7

Adidas-

Accounts receivable turnover = 10,251 / 1,530 = 6.7